What is the payback period for a project with the following costs: Initial investment: $46,700 Year 1 cash inflow: $10,000 Year 2 cash inflow: $10,000 Year 3 cash inflow: $12,000 Year 4 cash inflow: $12,000 Year 5 cash inflow: $12,000

3.23 years

3.41 years

3.79 years

4.23 years
Answer #1

012 - Fax +4+2700/12000 A B OS 2 year Cash flows Cumulative Cash flows 0 $ (46,700.00) $ (46,700.00) 4 1 $ 10,000.00 $ (36,70

012 fx =4+2700/12000 A 2 year 30 Cash flows -46700 10000 10000 12000 12000 12000 Cumulative Cash flows =B3 =C3+B4 =C4+B5 =C5+